# Weighted Chernoff information and optimal loss exponent in context-sensitive hypothesis testing

### 09 Mar 2026 v1

@@ 4.4 Extension to $M$-ary hypothesis testing
and taking infimum over $\delta_{n}$ gives $w_{\min}L_{n,M}^{*}\leq L_{n,M}^{(w)*}\leq w_{\max}L_{n,M}^{*}$. Hence $\frac{1}{n}\ln L_{n,M}^{(w)*}$ and $\frac{1}{n}\ln L_{n,M}^{*}$ have the same limit $C_{M}^{\mathrm{w}}$.
5 Conclusion
We studied context-sensitive (weighted) simple hypothesis testing under a multiplicative weight function $\varphi$ and showed that the optimal total loss (sum of type-I and type-II context-sensitive losses) admits an exact single-letter logarithmic asymptotic with exponent given by the weighted Chernoff information $D_{C}^{\mathrm{w}}$. A key technical step is an exponential-family embedding of the weighted geometric mixtures $\varphi\,p^{\alpha}q^{1-\alpha}$, which identifies the optimal Chernoff parameter $\alpha^{*}$ through the log-normaliser and yields information-geometric identities (including weighted Bregman representations). We also established finite-$n$ concentration bounds for the tilted weighted log-likelihood via martingale inequalities and derived explicit closed-form formulas for $D^{\mathrm{w}}_{B,\alpha}$ and $D_{C}^{\mathrm{w}}$ in several standard parametric models (Gaussian, Poisson, exponential), recovering the classical unweighted expressions when $\varphi\equiv 1$; an extension to finitely many hypotheses follows by reduction to pairwise weighted Chernoff exponents.
6 Acknowledment
The work by MK was done in the framework of the Basic Research Program of HSE.
7 Appendix: Cauchy location–scale family
@@
In this appendix we set $\varphi\equiv 1$, so the weighted quantities from Section 2 coincide with their classical counterparts (e.g. $\rho^{\mathrm{w}}_{\alpha}=\rho_{\alpha}$, $D^{\mathrm{w}}_{B,\alpha}=D_{B,\alpha}$, $D_{C}^{\mathrm{w}}=D_{C}$). This example provides a closed-form illustration outside the exponential-family setting.
Although the main body of the paper develops weighted Chernoff/Bhattacharyya quantities, we include the Cauchy location–scale family as an unweighted baseline closed-form benchmark outside the exponential-family setting (note the appearance of complete elliptic integrals even in the classical case). This appendix is not used in the proofs of the weighted results; it serves as a sanity check and illustrates the analytic complexity of $\rho_{\alpha}$ beyond exponential families. Nontrivial weights must satisfy the finiteness conditions from Section 2; for heavy-tailed Cauchy laws, common exponential tilts $\varphi(x)=e^{\gamma x}$ violate these conditions for any $\gamma\neq 0$, and in general a nonconstant weight also breaks the symmetry leading to $\alpha^{*}=1/2$.
KL divergence (closed form).
Proposition 7.1 (KL divergence between two Cauchy laws).
@@
$D_{\rm KL}(\mathbb{P}||\mathbb{Q})=\int_{\mathbb{R}}p_{l_{1},s_{1}}(x)\,\ln\frac{p_{l_{1},s_{1}}(x)}{p_{l_{2},s_{2}}(x)}\,\mathrm{d}x=\ln\frac{(s_{1}+s_{2})^{2}+(l_{1}-l_{2})^{2}}{4s_{1}s_{2}}.$ (7.1)
Chernoff parameter and Bhattacharyya coefficient.
@@
Proposition 7.2 (Chernoff parameter for Cauchy).
@@
$\rho_{\alpha}(\mathbb{P},\mathbb{Q})=\rho_{1-\alpha}(\mathbb{P},\mathbb{Q}),\qquad\alpha\in[0,1],$ (7.2)
equivalently $D_{B,\alpha}(\mathbb{P},\mathbb{Q})=D_{B,1-\alpha}(\mathbb{P},\mathbb{Q})$. Consequently, if $\mathbb{P}\neq\mathbb{Q}$, then the Chernoff maximiser is unique and equals
@@
Set $F(\alpha):=\ln\rho_{\alpha}(\mathbb{P},\mathbb{Q})$. By Hölder’s inequality, $F$ is convex on $[0,1]$, and it is strictly convex unless $p_{l_{1},s_{1}}/p_{l_{2},s_{2}}$ is $\mu$-a.e. constant (equivalently, unless $\mathbb{P}=\mathbb{Q}$). Hence $D_{B,\alpha}(\mathbb{P},\mathbb{Q})=-F(\alpha)$ is concave (strictly concave when $\mathbb{P}\neq\mathbb{Q}$). For Cauchy laws, the symmetry (7.2) holds; see, e.g., [8] for an invariance-based proof. Therefore $D_{B,\alpha}$ is symmetric about $1/2$, and concavity implies that it attains its maximum at $\alpha=1/2$. Strict concavity yields uniqueness when $\mathbb{P}\neq\mathbb{Q}$. ∎
Closed form for $\rho_{1/2}$ and $D_{C}$.
Lemma 7.3 (A standard elliptic-integral identity).
@@
$\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{\sqrt{(x^{2}+a^{2})\bigl((x-d)^{2}+b^{2}\bigr)}}=\frac{4}{\sqrt{(a+b)^{2}+d^{2}}}\,K\!\left(\frac{(a-b)^{2}+d^{2}}{(a+b)^{2}+d^{2}}\right).$ (7.3)
Proposition 7.4 (Bhattacharyya coefficient for Cauchy, closed form).
@@
$\rho_{1/2}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}\sqrt{p_{l_{1},s_{1}}(x)\,p_{l_{2},s_{2}}(x)}\,\mathrm{d}x=\frac{4\sqrt{s_{1}s_{2}}}{\pi\sqrt{(s_{1}+s_{2})^{2}+\delta^{2}}}\,K\!\left(\frac{(s_{1}-s_{2})^{2}+\delta^{2}}{(s_{1}+s_{2})^{2}+\delta^{2}}\right).$ (7.4)
@@
$D_{C}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q})=-\ln\rho_{1/2}(\mathbb{P},\mathbb{Q}),$ (7.5)
with $\rho_{1/2}(\mathbb{P},\mathbb{Q})$ given by (7.4).
@@
Apply Lemma 7.3 with $a=s_{2}$, $b=s_{1}$, $d=\delta$ to get (7.4). Finally, Proposition 7.2 yields $D_{C}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q})$, hence (7.5). ∎
Remark 7.5 (Useful special cases).
If $l_{1}=l_{2}$ (common location), then $\delta=0$ and (7.4) reduces to
@@
Remark 7.6 (Weighted case).
For a general weight $\varphi$, one has $\rho^{\mathrm{w}}_{1/2}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}\varphi(x)\sqrt{p_{l_{1},s_{1}}(x)p_{l_{2},s_{2}}(x)}\,\mathrm{d}x$. The symmetry (7.2) (and hence $\alpha^{*}=1/2$) typically fails unless $\varphi$ is compatible with the invariance argument used in [8]. We therefore use the Cauchy model mainly as a closed-form baseline example at $\varphi\equiv 1$.

### 09 May 2026 v3

@@ 4.5 Extension to $M$-ary Hypothesis Testing
and taking infimum over $\delta_{n}$ gives $w_{\min}L_{n,M}^{*}\leq L_{n,M}^{(w)*}\leq w_{\max}L_{n,M}^{*}$. Hence, $\frac{1}{n}\ln L_{n,M}^{(w)*}$ and $\frac{1}{n}\ln L_{n,M}^{*}$ have the same limit $C_{M}^{\mathrm{w}}$.
5 Conclusions
We studied context-sensitive simple hypothesis testing under a multiplicative weight and proved that the optimal total loss admits the single-letter logarithmic asymptotic
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}.$
The rate is the weighted Chernoff information. The main structural ingredient is an exponential-family embedding of the weighted geometric mixtures $\varphi\,p^{\alpha}q^{1-\alpha}$, which yields the characterisation of the optimal Chernoff parameter through the log-normaliser and leads to weighted information-geometric identities. We also derived finite-$n$ concentration bounds for the tilted weighted log-likelihood, obtained explicit formulas in Gaussian, Poisson, and exponential models, and extended the logarithmic asymptotic to finitely many hypotheses through the minimum pairwise weighted Chernoff information.
Open Problems
The single-letter representation (4) rests on Assumption 1.1 (factorised weights), on the integrability of the tilted log-normaliser $\hat{F}$, and on i.i.d. sampling of simple hypotheses. Relaxing these assumptions defines several natural open directions.
  • (a)
    Non-factorised weights. Replacing Assumption 1.1 by a weight $\varphi(x_{1}^{n})=\psi\bigl(\tfrac{1}{n}\sum_{i=1}^{n}h(x_{i})\bigr)$ or a pairwise-interaction weight; the single-letter rate is then expected to be replaced by a variational formula over the space of probability measures, in the spirit of Sanov/Gibbs conditioning.
  • (b)
    Integrability. Weights with heavy tails in the sufficient statistic $t(x)$ may violate the finiteness of $\hat{F}$ near $\alpha^{\ast}\theta_{1}+(1-\alpha^{\ast})\theta_{2}$; the boundary cases $\alpha^{\ast}\in\{0,1\}$ (cf. Proposition 3.10 and [3]) call for a systematic treatment via truncation or a change in base measure.
  • (c)
    Dependent observations. A weighted counterpart of the Gärtner–Ellis theorem for stationary ergodic sequences, along the lines of the weighted extensions in [4].
  • (d)
    Composite hypotheses. A weighted analogue of the generalised likelihood-ratio test and its exponent, with sup/inf characterisations in terms of $D_{C}^{\mathrm{w}}$ over the composite parameter sets.
  • (e)
    Information geometry of weighted manifolds. Extending the Chentsov–Amari framework [9, 8, 10, 11, 12] to weighted statistical manifolds; the Fisher metric and the dually flat structure depend on symmetries that $\varphi$ breaks in a controlled way.
Author Contributions
Conceptualization, methodology, and supervision, M.K.; formal analysis and validation, M.K. and E.Y.K.; software, visualization, data curation, and project administration, E.Y.K.; writing–original draft preparation, M.K. and E.Y.K.; writing–review and editing, E.Y.K.; funding acquisition, M.K. and E.Y.K. All authors have read and agreed to the published version of the manuscript.
Funding
The work by MK was carried out in the framework of a research project HSE-BR-2025-039 implemented as part of the Basic Research Program at HSE University. The second author (E.Yu. Kalimulina) was supported by the Ministry of Science and Higher Education of the Russian Federation under the state assignment (project FFNU-2025-0029).
Data Availability Statement
The Python/Jupyter notebook reproducing Table 1 and Figures 1–3, together with the direct numerical-integration verification of the Gaussian formula in Section 4.1, is openly available on Zenodo. No new experimental datasets were generated.
Conflicts of Interest
The authors declare no conflicts of interest. The funders had no role in the design of the study; in the collection, analyses, or interpretation of data; in the writing of the manuscript; or in the decision to publish the results.
Appendix A Cauchy Location–Scale Family
@@
In this appendix, we set $\varphi\equiv 1$, so the weighted quantities from Section 2 coincide with their classical counterparts (e.g., $\rho^{\mathrm{w}}_{\alpha}=\rho_{\alpha}$, $D^{\mathrm{w}}_{B,\alpha}=D_{B,\alpha}$, $D_{C}^{\mathrm{w}}=D_{C}$). This example provides a closed-form illustration outside the exponential-family setting.
Although the main body of the paper develops weighted Chernoff/Bhattacharyya quantities, we include the Cauchy location–scale family as an unweighted baseline closed-form benchmark outside the exponential-family setting (note the appearance of complete elliptic integrals even in the classical case). This appendix is not used in the proofs of the weighted results; it serves as a sanity check and illustrates the analytic complexity of $\rho_{\alpha}$ beyond exponential families. Nontrivial weights must satisfy the finiteness conditions from Section 2; for heavy-tailed Cauchy laws, common exponential tilts $\varphi(x)=e^{\gamma x}$ violate these conditions for any $\gamma\neq 0$, and in general, a nonconstant weight also breaks the symmetry leading to $\alpha^{*}=1/2$.
A.1 KL Divergence (Closed Form)
Proposition A.1 (KL divergence between two Cauchy laws).
@@ A.1 KL Divergence (Closed Form)
$D_{\rm KL}(\mathbb{P}||\mathbb{Q})=\int_{\mathbb{R}}p_{l_{1},s_{1}}(x)\,\ln\frac{p_{l_{1},s_{1}}(x)}{p_{l_{2},s_{2}}(x)}\,\mathrm{d}x=\ln\frac{(s_{1}+s_{2})^{2}+(l_{1}-l_{2})^{2}}{4s_{1}s_{2}}.$ (67)
A.2 Chernoff Parameter and Bhattacharyya Coefficient
@@
Proposition A.2 (Chernoff parameter for Cauchy).
@@
$\rho_{\alpha}(\mathbb{P},\mathbb{Q})=\rho_{1-\alpha}(\mathbb{P},\mathbb{Q}),\qquad\alpha\in[0,1];$ (68)
equivalently, $D_{B,\alpha}(\mathbb{P},\mathbb{Q})=D_{B,1-\alpha}(\mathbb{P},\mathbb{Q})$. Consequently, if $\mathbb{P}\neq\mathbb{Q}$, then the Chernoff maximiser is unique and equals
@@
Set $F(\alpha):=\ln\rho_{\alpha}(\mathbb{P},\mathbb{Q})$. By Hölder’s inequality, $F$ is convex on $[0,1]$, and it is strictly convex unless $p_{l_{1},s_{1}}/p_{l_{2},s_{2}}$ is $\mu$-a.e. constant (equivalently, unless $\mathbb{P}=\mathbb{Q}$). Hence $D_{B,\alpha}(\mathbb{P},\mathbb{Q})=-F(\alpha)$ is concave (strictly concave when $\mathbb{P}\neq\mathbb{Q}$). For Cauchy laws, the symmetry (68) holds; see, e.g., [17] for an invariance-based proof. Therefore, $D_{B,\alpha}$ is symmetric about $1/2$, and concavity implies that it attains its maximum at $\alpha=1/2$. Strict concavity yields uniqueness when $\mathbb{P}\neq\mathbb{Q}$. ∎
A.3 Closed Form for $\rho_{1/2}$ and $D_{C}$
Lemma A.3 (A standard elliptic-integral identity).
@@ A.3 Closed Form for $\rho_{1/2}$ and $D_{C}$
$\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{\sqrt{(x^{2}+a^{2})\bigl((x-d)^{2}+b^{2}\bigr)}}=\frac{4}{\sqrt{(a+b)^{2}+d^{2}}}\,K\!\left(\frac{(a-b)^{2}+d^{2}}{(a+b)^{2}+d^{2}}\right).$ (69)
Proposition A.4 (Bhattacharyya coefficient for Cauchy, closed form).
@@
$\rho_{1/2}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}\sqrt{p_{l_{1},s_{1}}(x)\,p_{l_{2},s_{2}}(x)}\,\mathrm{d}x=\frac{4\sqrt{s_{1}s_{2}}}{\pi\sqrt{(s_{1}+s_{2})^{2}+\delta^{2}}}\,K\!\left(\frac{(s_{1}-s_{2})^{2}+\delta^{2}}{(s_{1}+s_{2})^{2}+\delta^{2}}\right).$ (70)
@@
$D_{C}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q})=-\ln\rho_{1/2}(\mathbb{P},\mathbb{Q}),$ (71)
with $\rho_{1/2}(\mathbb{P},\mathbb{Q})$ given by (70).
@@
Apply Lemma A.3 with $a=s_{2}$, $b=s_{1}$, $d=\delta$ to get (70). Finally, Proposition A.2 yields $D_{C}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q})$, hence (71). ∎
Remark A.5 (Useful special cases).
If $l_{1}=l_{2}$ (common location), then $\delta=0$ and (70) reduces to
@@
Remark A.6 (Weighted case).
For a general weight $\varphi$, one has
$\rho^{\mathrm{w}}_{1/2}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}\varphi(x)\sqrt{p_{l_{1},s_{1}}(x)p_{l_{2},s_{2}}(x)}\,\mathrm{d}x.$
The symmetry (68) (and hence $\alpha^{*}=1/2$) typically fails unless $\varphi$ is compatible with the invariance argument used in [17]. We therefore use the Cauchy model mainly as a closed-form baseline example at $\varphi\equiv 1$.